Given the F2 progeny of 814 and 788 parentals and 501 and 260 recombinant types what is the frequency of recombination correct to 2 decimal places?
Answer #1

Total population of the F2 generation is 814 + 788 + 501 + 260 = 2363

Recombinant population = 501 + 260 = 761

Frequency of recombination = 0.3220 = 0.32